# AdaMesh: Personalized Facial Expressions and Head Poses for Adaptive Speech-Driven 3D Facial Animation

### v3

@@ 5.3 Approach Analysis
Table 3: Ablation for pose adapter.
@@
Ours 0.90 0.87 1.57
w/o style matrix 1.01 0.69 0.69
w/o HuBERT feature 0.91 0.07 0.15
w/o VQ-VAE 0.91 0.05 0.02
@@
Figure 4: Approach analysis. (a) performance vs. amount of adaptation data. (b) performance vs. amount of trainable parameters. (c) yaw degree curves for head poses. (d) t-SNE visualization of the semantic-aware pose style matrices.
6 Conclusion
@@
Appendix A Appendix
This supplemental material contains three subsections: the details of the expression adapter and the pose adapter in AdaMesh, implementations of baseline methods, and additional information about the user study.
A.1 Details of expression adapter and pose adapter
The architecture details of AdaMesh are presented in Tab. 4. The overall architecture of the expression adapter is illustrated in Fig. 1 of the main paper. For the identity encoder, $T$ frames identity parameters are fed into two convolutional layers and average to a vector with a dimension of 256. The speech sequence and expression sequence are separately processed by the audio encoder and style encoder. These two encoders are three conformer layers to output features with $T\times 256$ dimension. The encoded features are concatenated with the identity latent feature and sent into the expression decoder. We also adopt three Conformer layers for the decoder to output the predicted expression sequence. This is a simple architecture, but with robust and efficient performance, which is mainly attributed to the identity encoder and the style encoder providing the residual style information.
For the pose adapter, following Siyao et al. (2022), VQ-VAE downsample and upsample the pose sequence with a symmetrical network structure of 1D convolutional layers. We also adopt the cross-conditional causal attention mechanism other than the vanilla attention in Transformer, since the head poses of different degrees are entangled with each other. We present the loss function of VQ-VAE as below:
$\mathcal{L}_{VQ}=\mathcal{L}_{rec}(\hat{\bm{\psi}},\bm{\psi})+\left\|sg[\hat{z}]-z_{q}\right\|+\beta\left\|z-sg[z_{q}]\right\|$ (4)
where $\hat{\bm{\psi}}$ and $\bm{\psi}$ denotes output euler angle sequence and input ground truth, respectively. And the specific formula for $\mathcal{L}_{\text{rec}}$ can be expressed as:
$\mathcal{L}_{\text{rec }}(\hat{\bm{\psi}},\bm{\psi})=\left\|\hat{\bm{\psi}}-\bm{\psi}\right\|_{1}+\alpha_{1}\left\|\hat{\bm{\psi}}^{\prime}-\bm{\psi}^{\prime}\right\|_{1}+\alpha_{2}\left\|\hat{\bm{\psi}}^{\prime\prime}-\bm{\psi}^{\prime\prime}\right\|_{1}$ (5)
where $\bm{\psi}^{\prime}$ and $\bm{\psi}^{\prime\prime}$ represent the 1st-order (velocity) and 2nd-order (acceleration) partial derivatives of euler angle sequence $\bm{\psi}$ on time. In addition, $\alpha_{1}$, $\alpha_{2}$ and $\beta$ denote trade-off coefficients while $sg[\cdot]$ denotes ”stop gradient”.
Table 4: Architecture parameter list for AdaMesh.
Expression Adapter Identity Encoder Conv1d Kernel 5
Conv1d Channel Size 256
Identity Embedding Width 256
Audio Encoder Conformer Layer 3
Conformer Layer Head 4
Conformer Layer Embedding Width 256
Audio Embedding Width 256
Style Encoder Conformer Layer 3
Conformer Layer Head 4
Conformer Layer Embedding Width 256
Audio Embedding Width 256
Expression Decoder Conformer Layer 3
Conformer Layer Head 4
Conformer Layer Embedding Width 256
Audio Embedding Width 256
Conv1d Kernel 5
Conv1d Channel Size 256
Pose Adaptor VQ-VAE Encoder Layers 3
Decoder Layers 3
Encoder/Decoder Conv1d Kernel 4
Encoder/Decoder Conv1d Channel Size 256
Codebook Size 1024
Codebook Embedding Width 256
Pose GPT Transformer Layers 6
Transformer Layer Head 12
Transformer Layer Embedding Width 768
Style Embedding Width 50
A.2 Implementations of baseline methods
As mentioned in the main paper, we compare AdaMesh with state-of-the-art methods. FaceFormer (Fan et al., 2022) and CodeTalker (Xing et al., 2023) have released the pre-trained model, thus we test inference with the available codes. To calculate metrics with the same identity topology, we adopt Li et al. (2017) to recover the shape parameter from 2D video are convert it to the format that FaceFormer and CodeTalker require for inference. For GeneFace (Ye et al., 2023), we extract the required data to train the image render and fine-tune the post-net. For FACIAL (Zhang et al., 2021), since it is a person-specific method which requires a large amount of data for training, we directly utilize the pre-trained checkpoint of Obama for inference. For Imitator (Thambiraja et al., 2023), it is the most-related and latest work. We implement it to the best of our understanding. As soon as Imitator releases the public codes, we will replace the visualized images in our paper with the original ones.
A.3 User Study
The designed user study of the mean-of-score tests for lip synchronization and head pose naturalness and the A/B tests for the richness of facial expressions are separately demonstrated in Fig. 5 and Fig. 6. Prior to commencing the tests, we have thoroughly communicated the experimental guidelines and important information that users need to be aware of. For the lip synchronization and head pose naturalness tests, we sample 105 videos for seven compared methods. For the A/B tests, 60 pairs of compared videos are created.
Figure 5: Designed mean-of-score test interface for lip synchronization and head pose naturalness.
Figure 6: Designed A/B test interface for the richness of facial expressions.

### v4

@@ V-B Quantitative Evaluation
TABLE IV: Ablation study for pose adapter.
@@
Ours 0.90 0.87 1.57 13.89
w/o style matrix 1.01 0.69 0.69 19.88
w/o HuBERT feature 0.91 0.07 0.15 26.73
w/o VQ-VAE 0.91 0.05 0.02 22
Comparison on VOCASET. Our approach is designed for personalized facial expressions and head poses. However, VOCASET [1] has no explicit talking styles or head poses. To evaluate the generalization ability of our approach, we convert the VOCASET vertices to expression coefficients for inference. We test our approach on two settings. 1) Ours w/o FT: we directly run the inference pipeline of the expression adapter with the unseen speech in VOCASET without fine-tuning any parameters. The converted expressions are sent into the style encoder of the expression adapter. 2) Ours w/ FT: we fintune the MoLoRA parameters of the expression adapter with the converted expressions and speech data. Since the converted expressions are not as accurate as the original vertices, we train FaceFormer [3] and CodeTalker [4] on the converted data for a fair comparison. The lip vertex error (LVE) calculated in Tab.II indicates that our approach achieves competitive performance, although AdaMesh is not directly trained on VOCASET. This confirms the generalization ability of our approach. When the MoLoRA parameters in the expression adapter are updated with only 30 seconds of data, our approach gains better performance than other methods.
Figure 6: Expression adapter analysis. (a) Performance vs data amount. (b) Performance vs adaptation parameter amount. Lower L2 error of vertices and higher diversity score of expressions indicate better performance. The settings with the best performance are annotated with the yellow crowns.
TABLE V: Performance of MoLoRA and LoRA with different rank size settings. Values with the same single or wavy underlines have approximately the same number of trainable parameters and can be compared.
[…]
MoLoRA [4,8,16,32] 0.7M 2.25 2.91 7.10
[4,32] 0.4M 2.32 3.07 6.52
[8,16] 0.3M 2.36 3.10 6.21
LoRA 128 1.5M 2.29 3.00 6.34
64 0.7M 2.38 3.11 5.80
36 0.4M 2.42 3.07 5.95
V-C Approach Analysis
Our approach involves multiple modules and training settings, thus we conduct extensive analysis to show the effectiveness of these configures.
Ablation Study for Expression Adapter. For expression adapter, we separately remove the following modules to verify their necessity. 1) As shown in Tab. III, the removal of the style encoder leads to a drop of vertices errors and diversity metrics by a large margin. It suggests that the style encoder provides essential details which cannot be predicted from speech. 2) The absence of identity also causes a slight performance drop since the expression and identity parameters are not fully disentangled [19, 38]. The identity parameters also contain person-specific talking styles. 3) We fine-tune the full expression adapter instead of updating MoLoRA parameters. It can be observed that the performance significantly decreases since fine-tuning with few data damages the generalization ability of the pre-trained model.
Ablation Study for Pose Adapter. For pose adapter, we test three settings in Tab.IV. 1) The retrieval of the pose style matrix is removed by randomly selecting a pose style embedding in the training data for inference. All metrics drop significantly due to the unawareness of the reference’s personalized talking style. 2) w/o HuBERT feature, where only the mel spectrogram is utilized for the speech representation. The LSD and FSD scores show that HuBERT features contain semantic information, which is vital to the diverse and meaningful head pose generation. 3) After removing VQ-VAE, the model directly predicts the continuous poses. The diversity scores drop to near zero. It suggests that VQ-VAE provides realistic quantized codes which is easier to predict. To further investigate the efficacy, we plot the yaw degree of head poses in Fig. 7(a). The pose curve of AdaMesh has the largest variation, which is closer to the ground truth. It confirms the results in quantitative and qualitative experiments. We also utilize t-SNE [51] to visualize the pose style matrix for the selected exciting and peaceful head poses. The visualization result shows that different styles are clustered into different groups. This proves that the semantic-aware pose style matrix effectively aggregates and compresses different pose styles, thereby making the representation of pose styles not isolated and discrete, but continuous and meaningful. A similar example is that the inputs of BERT [52] are one-hot discrete word embeddings, but the model learns meaningful representations with context-sensitive semantics.
@@ V-C Approach Analysis
Figure 7: Pose adapter analysis. (a) Yaw degree curves for head poses. (b) t-SNE visualization of the semantic-aware pose style matrices.
Figure 8: Correlation of speech features and head pose. Yaw degree is plotted as the head pose.
Amount of Adaptation Data. We only analyze the impact of the adaptation data amounts on the expression adapter performance, since the pose adapter actually is a zero-shot model. As shown in Fig. 6(a), points within the same color family represent the same method, while the same marker corresponds to the computation of the same metric. Thus the points with the same marker can be compared. The figure has two y-axises, and the legend next to each y-axis indicates what metric this y-axis represents. It can conclude that increasing the adaptation data amount can improve performance. AdaMesh achieves significantly better performance than Imitator with the same amount of adaptation data.
Figure 9: Visualization of head pose styles for different speakers. Description, one of the head pose degrees and the corresponding video are demonstrated.
Amount of Adaption Parameters. We estimate the impact of MoLoRA parameter amounts on the expression adapter performance. MoLoRA with different rank size combinations and the vanilla LoRA (the point in Fig. 6(b) with the maximum parameter amount) are compared. The point with the highest diversity score corresponds to the rank size combination of $[4,8,16,32]$. This setting also achieves the lowest LVE and EVE scores. Furthermore, when optimizing roughly the same number of parameters, our approach outperforms Imitator in all metrics.
Analysis on MoLoRA Rank Size. Tab. V shows the performance of the expression adapter with different rank sizes. It shows that increasing the rank size tends to yield better performance. The rank size combination of [4,8,16,32] for MoLoRA achieves the best performance. We compare with the vanilla LoRA [15] of different rank sizes. It is worth noting that with the same number of trainable parameters, MoLoRA achieves better results than the vanilla LoRA, even surpassing the LoRA with the rank size of 128. This indicates that the design of MoLoRA, which combines LoRAs of different rank sizes, enhance model performance more effectively than increasing the number of trainable parameters.
Correlation between Speech and Head Poses. Due to the subtle nature of head movements, it’s challenging to observe their semantic correlations, especially when represented in a mesh format. Thus, we plot the speech features and the head pose curves in Fig. 8. The predicted and ground truth head poses are different in absolute value due to the one-to-many mapping between speech and pose. It is noticeable that significant changes in head poses correspond with variations in pitch and energy in speech (representing shifts in emphasis or emotion), demonstrating a positive correlation. Head poses distinctly change when the speech content involves emphasis or negation, indicating that the generated head poses contain semantic information, which is consistent with previous works [27, 28, 29]. The significant pose curve fluctuations on spoken words for the predicted head pose and the ground truth validates the reenactment of the reference style. Additionally, unstable head movements in the training dataset do not affect the generation of head poses, since the VQ-VAE, trained on a large dataset, effectively filtered out these unstable jitters. This can be confirmed by the user study.
Visualization of Specific Head Pose Styles. As shown in Fig. 9, we conduct a clustering analysis similar to Fig. 7(b) on the database for all pose style matrices and visualize some of the cluster centers and outliers. It demonstrates that different individuals have different head pose styles. By observing the corresponding ground-truth video and audio, we find that the head pose styles are influenced by multiple factors such as the scenes in which the speakers are situated, their personalities, and moods. This verifies the rationale for the targeted modeling of head pose styles.
Figure 10: Visual comparison with methods of different topologies. Emotion can be observed on AdaMesh results.
Figure 11: Qualitative comparison with EMO [53] on the portrait-realistic talking face generation task. English and Chinese results are given separately. We annotate the closest equivalent English phonetic symbols next to the Chinese pronunciation for better observation.
V-D More Demonstration
Comparison with Other Topologies. In Fig. 10, AdaMesh is compared with methods of different topologies [2, 22] that cannot be easily converted to FLAME format [19]. AdaMesh still achieves better lip sync and richer facial expressions.
Portrait-Realistic Talking Face Generation. In Fig.11, we further show that our model can generate coherent photo-realistic video by employing a simple conditional GAN-based person-specific image renderer [23, 24]. Expressions and head poses are predicted by the proposed AdaMesh. Then the image renderer takes the rasterized 2D mesh sequence and the identity image sequence as inputs and synthesizes a portrait video, which matches the expression details of the mesh sequence and the appearance of the identity sequence. We provide the results of a state-of-the-art talking face method EMO [53] as a comparison to AdaMesh. EMO is an end-to-end method that directly predicts image variation from speech without any intermediate face representation, thus achieving highly synchronized lip movements. AdaMesh gains comparable lip synchronization performance with EMO. This suggests that our approach can predict accurate-enough expressions compared with the end-to-end method and can also serve as a conduit for photo-realistic avatars.
Generation Efficiency. When inference on Nvidia A10 GPU, it takes 10ms from inputting speech, predicting facial parameters, to generating meshes. Utilizing the renderer to synthesize portrait-realistic images will cost 9ms. Our method can be employed in the realistic human live-streaming scenario with a rapid inference speed at 48fps.
VI Conclusion